OpenStudy (anonymous):

which ordered pair is NOT a solution of the inequaality, 3x-2y>or equal to 12. Please help! a. (-3, -10.5) b.(-1,-7) c. (o,-6) d. (2,-3)

OpenStudy (anonymous):

put each of them into the formula -3*3-2*-10.5=12 (A is solution) -1*3-2*-7=11 (B is not solution) 0*3-2*-6=12 (C is solution) 2*3-2*-3=12 (D is solution)